A rice dealer mixes two varieties of rice, one that sells at P500 a sack and anotherat P520 a sack. He intends to fill up 100 sacks to be sold at P515 each. How many sacks of rice each variety are needed?
Answer by mananth(16949)

let the number of P500 sacks he mixes be x
the P520 he mixes will be 100-x
..
mixture P515 100 bags
500x+520(100-x)=515*100
500x+52000-520x=51500
-20x=51500-52000
-20x=-500
x=-500/-20
x=25 -------- the p500 type
balance 75 will be P520 type sacks